If someone invests $3000 at 5% annual interest and $4000 at 6% annual interest, how much interest do they earn in one year? How many total dollars will they have at the end of the year?
Answer by stanbon(48551)

If someone invests $3000 at 5% annual interest and $4000 at 6% annual interest, how much interest do they earn in one year? How many total dollars will they have at the end of the year?
-----------------
5% investment DATA:
Amount = 3000 ; Interest = 0.05(3000)= 150 dollars
-------
6% investment DATA:
Amount = 4000 ; Interest = 0.06(4000)= 240 dollars
--------
Total interest = interest + interest
Tatal = 150 + 240 = $390